Ohme Framing's cost formula for its supplies cost is $1,620 per month plus $13 per frame. For the month of April, the company planned for activity of 882 frames, but the actual level of activity was 878 frames. The actual supplies cost for the month was $13,500. The supplies cost in the flexible budget for April would be closest to:
Answer #1
Concepts and reason

Budget: The budget is a plan which is prepared for a specific period for expenditures or incomes related to business. The objective to prepare the budget is to compare the estimates belong to specific expenditure or income with the actual figures for the purpose of controlling the cost and reducing the cost.

Fundamentals

Flexible budget: The budget which is prepared to control the day-to-day operations of the business. To make the control effective, an estimated statement is prepared and compared with the report of actual performance. The budget which classify the difference between fixed cost, semi-variable cost and variable cost which changes with the level of activity performed is flexible budget.

Variable cost: The cost and expenses which change with the level of activity. The expenses which change with every extra unit produced. More the company produce, more will be the variable cost.

Fixed Cost: It is the cost which remains the same irrespective of the level of output. The fixed cost remains unaffected by change in the level of output. The per unit fixed cost reduces with the increase in the level of output and the fixed cost per unit increased with a decrease in units of output.

Compute the total variable cost for the month of April.

Variablecost=Actualnumberofframes×Costperframe=878frames×$13perframes=$11,414\begin{array}{c}\\{\rm{Variable}}\,{\rm{cost}} = {\rm{Actual}}\,{\rm{number}}\,{\rm{of}}\,{\rm{frames}} \times {\rm{Cost}}\,{\rm{per}}\,{\rm{frame}}\\\\ = 878\,{\rm{frames}} \times \$ 13\,{\rm{per}}\,{\rm{frames}}\\\\ = \$ 11,414\\\end{array}

Compute the budgeted cost for the month of April.

BudgetedcostofApril=Fixedcost+Variablecost=$1,620+$11,414=$13,034\begin{array}{c}\\{\rm{Budgeted}}\,{\rm{cost}}\,{\rm{of}}\,{\rm{April}} = {\rm{Fixed}}\,{\rm{cost}} + {\rm{Variable}}\,{\rm{cost}}\\\\ = \$ 1,620 + \$ 11,414\\\\ = \$ 13,034\\\end{array}

Ans:

The flexible budgeted cost for the month of April is $13,034.